40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 9 20:01:25 2019, Jon, Update, VAC, Second pumpdown completed Screenshot_from_2019-01-09_20-00-39.pngIMG_3088.jpg
    Reply  Thu Jan 10 10:23:46 2019, gautam, Update, VAC, overnight leak rate OvernightLeak.png
       Reply  Sat Feb 2 21:53:24 2019, Koji, Update, VAC, overnight leak rate Screen_Shot_2019-02-02_at_21.49.33.png
          Reply  Sun Feb 3 15:15:21 2019, gautam, Update, VAC, overnight leak rate Jan10_data.pngFeb1_data.png
             Reply  Sun Feb 3 20:52:34 2019, Koji, Update, VAC, overnight leak rate 
                Reply  Mon Feb 4 12:23:24 2019, gautam, Update, VAC, pumpdown 83 - leak tests PD83.png
                   Reply  Tue Feb 5 10:11:30 2019, gautam, Update, VAC, leak tests complete, pumpdown 83 resumed Annuli.pngMainVol.png
                      Reply  Tue Feb 5 19:30:14 2019, gautam, Update, VAC, Main volume at 20 uTorr PD83.png
                         Reply  Thu Feb 7 13:55:25 2019, gautam, Update, VAC, RGA turned on 
                            Reply  Thu Feb 7 19:28:46 2019, gautam, Update, VAC, IFO recovery 
Message ID: 14438     Entry time: Thu Feb 7 13:55:25 2019     In reply to: 14436     Reply to this: 14440
Author: gautam 
Type: Update 
Category: VAC 
Subject: RGA turned on 

[chub, steve, gautam]

Steve came by the lab today. He advised us to turn the RGA on again, now that the main volume pressure is < 20 uTorr. I did this by running the RGAset.py script on c0rga - the temperature of the unit was 22C in the morning, after ~3 hours of the filament being turned on, the temperature has already risen to 34 C. Steve says this is normal. We also opened VM1 (I had to edit the interlocks.yaml to allow VM1 to open when CC1 < 20uTorr instead of 10uTorr), so that the RGA volume is exposed to the main volume. So the nightly scans should run now, Steve suggests ignoring the first few while the pumpdown is still reaching nominal pressure. Note that we probably want to migrate all the RGA stuff to the new c1vac machine.

Other notes from Steve:

  • RP1 and RP3 should have their oil fully changed (as opposed to just topped up)
  • VABSSCI adn VABSSCO are NOT vent valves, they are isolating the annuli of the IOO and OMC chambers from the BS chamber annuli. So next time we vent, we should fix this!
  • Leak rate of 3-5 mTorr/day is "normal" once the system has been pumped for a few days. Steve agrees that our observations of the main volume pressure increase is expected, given that we were at atmosphere.
  • Regarding the upcoming CES construction
    • Steve recommends keeping the door along the east arm, as it is useful for bringing equipment into the lab (end door access is limited because of end optical tables)
    • Particle counter data logging should be resumed before the construction starts, so that we can monitor if the lab is getting dirtier
  • OSEM filters (new ones, i.e. made according to spects in D000209) are in the Clean Cabinet (EX). They are individually packaged in little capsules, see Attachment #1. So the ones I installed were actually a 2002 vintage. We have 50pcs, enough to install new ones on all the core optics + spares.
ELOG V3.1.3-